### v3

# JaxRobotarium: Training and Deploying Multi-Robot Policies in 10 Minutes

Abstract
Multi-agent reinforcement learning (MARL) has emerged as a promising solution for learning complex and scalable coordination behaviors in multi-robot systems. However, established MARL platforms (e.g., SMAC and MPE) lack robotics relevance and hardware deployment, leaving multi-robot learning researchers to develop bespoke environments and hardware testbeds dedicated to the development and evaluation of their individual contributions. The Multi-Agent RL Benchmark and Learning Environment for the Robotarium (MARBLER) is an exciting recent step in providing a standardized robotics-relevant platform for MARL, by bridging the Robotarium testbed with existing MARL software infrastructure. However, MARBLER lacks support for parallelization and GPU/TPU execution, making the platform prohibitively slow compared to modern MARL environments and hindering adoption. We contribute JaxRobotarium, a Jax-powered end-to-end simulation, learning, deployment, and benchmarking platform for the Robotarium. JaxRobotarium enables rapid training and deployment of multi-robot reinforcement learning (MRRL) policies with realistic robot dynamics and safety constraints, supporting both parallelization and hardware acceleration. Our generalizable learning interface provides an easy-to-use integration with SOTA MARL libraries (e.g., JaxMARL). In addition, JaxRobotarium includes eight standardized coordination scenarios, including four novel scenarios that bring established MARL benchmark tasks (e.g., RWARE and Level-Based Foraging) to a realistic robotics setting. We demonstrate that JaxRobotarium retains high simulation fidelity while achieving dramatic speedups over baseline (20x in training and 150x in simulation), and provides an open-access sim2real evaluation pipeline through the Robotarium testbed, accelerating and democratizing access to multi-robot learning research and evaluation. Our code is publicly available on github.
Keywords: Multi-Robot Learning, Sim2Real Deployment, Benchmarking, Jax
1 Introduction
Traditional approaches to multi-robot coordination have a rich history and offer robust solutions to standardized problems across a variety of domains [1, 2, 3, 4]. However, these approaches often rely on explicit specification of generalizable coordination mechanisms, which either requires considerable expertise or is virtually impossible in complex unstructured domains. Learning-based techniques, such as multi-robot reinforcement learning (MRRL), are emerging as promising alternatives that can offer a generalizable recipe to enable effective multi-robot coordination across various applications (e.g., scheduling [5], warehouse automation [6], and autonomous driving [7, 8]).
Despite its potential, learning-based algorithms are seldom deployed and evaluated on physical multi-robot platforms. Even when physical deployments are undertaken, they tend to rely on bespoke and closed platforms and testbeds, limiting reproducibility and benchmarking. In stark contrast, the multi-agent reinforcement learning (MARL) research community outside of robotics benefits from standardized and structured benchmarks, simulation, and learning environments (e.g., Multi-Agent Particle Environment (MPE) [9], StarCraft Multi-Agent Challenge (SMAC) [10], and JaxMARL [11]). However, these environments often abstract away robotics-relevant characteristics (e.g., robot dynamics and collision avoidance) and result in a considerable sim2real gap [8], limiting their ability to faithfully evaluate learning approaches tailored to multi-robot systems.
MARBLER [12] is a recent framework that attempts to meet this critical need. MARBLER acts as a bridge between the standardized infrastructure designed for MARL algorithms and the Robotarium [13] (an open and remotely-accessible multi-robot testbed). However, MARBLER does not support parallelization and execution on GPU/TPU, resulting in prohibitively slow training speeds that hinder adoption and comprehensive evaluations.
In this work, we contribute JaxRobotarium, an end-to-end open-source Jax-based platform designed to significantly accelerate the training and deployment of MRRL algorithms while maintaining ease-of-use and open access to hardware-based evaluations (see Fig. 1).
JaxRobotarium includes three key contributions. First, we contribute a new Jax-based simulator for the Robotarium that is designed from the ground-up and modeled after the existing Robotarium Python Simulator (RPS). Our new simulator supports parallelization and GPU/TPU execution, enabling dramatic improvements in efficiency by reducing the time and computational overhead typically associated with learning multi-robot coordination in realistic settings. Second, we contribute a new intuitive interface to facilitate integration with existing MARL libraries, such as JaxMARL [11], allowing researchers to leverage and build upon existing algorithmic implementations with minimal setup. Third, we contribute a new unified benchmark for MRRL that currently supports 8 representative multi-robot coordination scenarios, with flexible tools to rapidly create new ones or customize the existing ones. The unique benefits of JaxRobotarium include:
Dramatic improvements in efficiency of training (up to 20x) and simulating (up to 150x) MRRL policies, supporting parallelization and execution on GPU/TPUs.
Support for realistic simulation of robots, adhering to their dynamic constraints and safety guarantees obtained from barrier certificates.
A standardized set of multi-robot coordination scenarios implemented in Jax with real-world counterparts for rapid benchmarking.
An open-source, no-cost platform that provides any user in the world with access to training and real-world deployment of MRRL algorithms.
We validate the utility and benefits of JaxRobotarium through extensive efficiency evaluations, algorithm benchmarking, and comprehensive sim2real transfer experiments. We demonstrate that, compared to MARBLER [12], JaxRobotarium simulates robot trajectories up to 150 times faster and trains multi-robot policies up to 20 times faster. We benchmarked four MARL algorithms on the 8 scenarios provided by JaxRobotarium, revealing new insights into the algorithms’ utility across different multi-robot coordination scenarios. We systematically studied the sim2real gap of JaxRobotarium using data from over 200 real-world deployments of multi-robot policies that were entirely trained in simulation. We note that JaxRobotarium is not a panacea for the sim2real gap. Rather, JaxRobotarium helps study and reduce the sim2real gap by enabling rapid training and deployment. Our analysis reveals that while many task-algorithm combinations result in negligible performance differences when deployed on hardware, certain combinations of MARL algorithms and task characteristics exhibit noticeable performance differences. However, we show that such differences can be largely mitigated through simple randomization techniques such as noisy actions. We believe JaxRobotarium will serve as a powerful tool for the multi-robot systems community to advance the development, benchmarking, and sim2real transfer of learning algorithms.
2 Related Work
The platforms to facilitate learning multi-robot coordination strategies should have the following characteristics: (i) support efficient training and evaluation of policies [11, 14]; (ii) have realistic simulations of robots [8, 13], taking into account their dynamic constraints and the need for lower-level collision avoidance; (iii) have a convenient interface with common MARL algorithms and benchmarking tasks [11, 12, 15]. Centering around these key requirements, we evaluate and compare available MRRL and MARL frameworks in Table 1. The Multi-Agent Particle Environment (MPE) [9] is a widely used framework for evaluating MARL algorithms on a diverse suite of 2D tasks. While MPE was enhanced by successors including VMAS [14], BenchMARL [15], and JaxMARL [11], none of them realistically reflects the dynamics of real robots, impeding straightforward sim2real evaluation of MRRL policies. IsaacGym [16] and IsaacLab [17] offer high-fidelity simulation but are geared toward single-robot learning across varied embodiments. Multi-robot learning in these frameworks requires additional infrastructure for integration with MARL libraries, collision avoidance, benchmark tasks, and a sim2real pipeline, none of which are natively provided. Thus, to better evaluate MRRL policies, researchers must build task-specific test platforms [18, 19, 20] or reproduce previously proposed test beds, such as Cambridge RoboMaster [21]. As a contrastive line of efforts, the Robotarium [13] is a publicly available, multi-robot testbed that supports testing the coordination of up to 20 GRITSBots [22] robots. The Robotarium provides a simulator with implemented collision avoidance using control barrier functions [23] and offers sim2real evaluation through its hardware test bed. It is the only publicly available and free-to-use multi-robot hardware testbed in the world, and has been widely used for multi-robot research in areas including task allocation [24], heterogeneous coordination [25], motion planning [26], and safety [27]. To bridge the gap between MARL and a real multi-robot testbed, [12] contributes MARBLER, which is an open platform that promotes the comprehensive evaluation of MRRL algorithms on the Robotarium. Unfortunately, MARBLER only runs on the CPU without parallelization, resulting in prolonged training and tedious development cycles. JaxRobotarium builds upon the philosophy of MARBLER, marrying it with the computational efficiency offered by Jax and GPU parallelization.
| Platform | Train On | Robot | Safety | Sim2Real | MARL | Open Access |
| GPU/TPU | Dynamics | Guarantee | Capabilities | Support | Hardware | |
| MARL Platforms (MPE [9], SMAC [10], VMAS [14], | ✓ | ✗ | ✗ | ✗ | ✓ | N/A |
| BenchMARL [15], JaxMARL [11], Google Research Football [28]) | ||||||
| IsaacGym [16], IsaacLab [17] | ✓ | ✓ | ✗ | ✓ | ✗ | ✗ |
| Cambridge RoboMaster [21] | ✓ | ✓ | ✗ | ✓ | ✓ | ✗ |
| MuSHR [19] | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ |
| Duckietown [20, 8] | ✓ | ✓ | ✗ | ✓ | ✓ | ✗ |
| Robotarium [13] | ✗ | ✓ | ✓ | ✓ | ✗ | ✓ |
| MARBLER [12] | ✗ | ✓ | ✓ | ✓ | ✓ | ✓ |
| JaxRobotarium (ours) | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ |
3 JaxRobotarium Platform
JaxRobotarium is designed to achieve the training speeds of Multi-Agent RL in Multi-Robot RL while offering standardized, public, and free sim2real evaluation through the Robotarium. To this end, JaxRobotarium consists of the following key components,
1
A Jax-based simulator for the Robotarium, compatible with parallelization and GPU/TPU execution.
2
A platform for Multi-Robot RL, which we integrated with JaxMARL [11] to enable training with several SOTA MARL algorithms.
3
An implementation of all four MARBLER tasks, and four newly implemented tasks.
4
A direct deployment pipeline for sim2real evaluation on the Robotarium.
We provide a full overview of the architecture in the Appendix (Figure 5).
4 Jax-based Robotarium Simulator
4.1 Design
We contribute Jax-RPS, a Jax simulator for the Robotarium. Jax-RPS uses a unicycle dynamics model to faithfully simulate the behavior of the Robotarium’s GRITSBots [29]. To control the GRITSBots, Jax-RPS implements single integrator and unicycle controllers to drive robots to desired positions/poses. To enforce the safety constraints often required in real-world deployment, Jax-RPS implements control barrier certificates for collision avoidance, which solve the quadratic program described in [30]. Finally, all functionality in Jax-RPS is designed to be compatible with Jax’s jit compilation for accelerated execution and vmap for parallelization. Overall, this results in Jax-RPS being optimized for data-driven paradigms, such as Multi-Agent RL, while maintaining its faithfulness to real-world conditions and deployment considerations.
4.2 Evaluation
To validate the faithfulness and optimization of Jax-RPS, we construct the Random Waypoint Navigation scenario where four robots must each navigate to a randomly selected waypoint before being assigned a new one. We conduct the following experiments:
1
We measure faithfulness by simulating Random Waypoint Navigation for 10K timesteps and computing the difference in robot trajectories between Jax-RPS and RPS across unicycle position and pose controllers. We report mean wall time and trajectory mismatch with associated standard error averaged across 30 trials in Table 2.
2
We measure optimization performance by simulating Random Waypoint Navigation for 100K timesteps, where the timesteps are collected across multiple environments in parallel. We modified RPS to support multi-processing for parallelization. We report mean wall time across the number of environments with associated standard deviation averaged over 30 trials in Figure 2.
In all trials, we randomize the order of the controller-simulator conditions to avoid ordering biases. The above experiments were run on an M4 Pro CPU and 24 GB of RAM.
As seen in Table 2, Jax-RPS simulates trajectories that are identical to RPS with a 140- 150x speedup over RPS when run on CPU. Furthermore, in Figure 2 we observe Jax-RPS continues to benefit from increased parallelization, with its wall time to collect 100K timesteps monotonically decreasing as the number of environments increases. In contrast, RPS suffers slightly at higher degrees of parallelization, where the overhead of managing the multiple processes outweighs the benefit of executing the environments in parallel.
| Controller | Simulator | Wall Time | Trajectory |
| (ms ) | Error () | ||
| Position | Default | 1425.28 4.79 | N/A |
| JaxRobotarium | 9.18 0.10 | 0.00 0.00 | |
| Pose | Default | 1713.87 5.39 | N/A |
| JaxRobotarium | 11.42 0.07 | 0.00 0.00 |
5 Benchmark Scenarios
We provide a diverse set of tasks to train on within JaxRobotarium. First, we re-implement the original MARBLER tasks in our framework, as described below.
Arctic Transport: Two drone robots that observe their surrounding terrain must help guide two traversal robots, one faster on ice and slower on water, and vice versa. The team is successful if both traversal robots reach the goal zone.
Predator Capture Prey Discovery: Sensing robots that sense landmarks within the environment collaborate with tagging robots that tag landmarks within the environment. Task performance is measured by the number of landmarks tagged. We note that this task has been renamed to Discovery for clarity.
Material Transport: Robots with varying speeds and capacities must collaborate to unload two material depots, one further with less material, and one closer with more material. Task performance is measured by the amount of material remaining to be unloaded.
Warehouse: Robots assigned to a zone type must collectively maximize the number of deliveries between their assigned zone type. Task performance is measured by the number of deliveries
In addition to the re-implementations above, we designed and implemented four new tasks:
Navigation (MAPF): Robots must navigate to independent goals within the environment. The team is successful if all robots are on their goals by the end of the rollout.
Foraging: Inspired by grid world Level-Based Foraging [31], robots with varying levels of foraging ability must collaborate to forage all resources in the environment. Task performance is measured by the number of resources foraged.
Continuous-RWARE: Inspired by the grid world RWARE [32], robots must collaborate to fulfill dropoffs of the requested shelves. Task performance is measured by the number of requested shelves dropped off.
Predator-Prey: Robots must collaborate to tag a more agile prey, where the prey is controlled by the heuristic proposed in [33]. Task performance is measured by the number of tags on the prey.
Visualizations of all scenarios can be seen in Figure 4, and detailed descriptions are in Appendix B.
6 Jax-Robotarium Interface
6.1 Design
To enable the training and deployment of multi-robot policies with our simulator, we build a learning interface consisting of 5 main components:
RobotariumEnv is the main interface structuring input to and output from Jax-RPS. It maintains a generalizable representation of the environment state for the Robotarium and provides visualization utilities to construct a consistent base visualization across all tasks.
HetMananger provides an interface for injecting heterogeneity into multi-robot environments. It supports common heterogeneous representations: robot IDs, class IDs, capability vectors, and can be extended to include more. It manages the robot’s awareness of heterogeneity, where the observation space can be configured to include any representation.
ControllerManager handles translating policy actions into commands to the robots. If a policy is designed to output waypoints, these waypoints are processed by the configured controller into unicycle velocity commands. If barrier functions are enabled, the unicycle velocity commands are then modified to guarantee collision avoidance.
Deploy is a module that streamlines deployment to the Robotarium by compiling all the necessary files for an experiment into a folder, which can then be directly uploaded and run on the Robotarium test bed. Since the Robotarium test bed does not support Jax, the deployment module performs model conversion and file sanitation to ensure compatibility.
We choose JaxMARL as the MARL framework, given its SOTA performance and extensive baseline algorithm implementations. However, note that any Jax multi-agent learning framework designed for the output structure of the Scenario interface is compatible with JaxRobotarium.
6.2 Evaluation
To validate the utility of our framework, we compare the training performance and efficiency of JaxRobotarium with no parallelization (1 env) and parallelization (8 envs) against MARBLER on the four MARBLER scenarios: Arctic Transport, Discovery, Material Transport, and Warehouse. We report the episodic returns and timesteps simulated against the wall time of the fastest run, averaged across 3 seeds. All experiments are run on an Intel i7-12700KF and NVIDIA RTX 3070.
In Figure 3, we find that JaxRobotarium can train for up to 20x more timesteps and achieve consistently higher returns compared to MARBLER given the same amount of time. As expected, the support for GPU execution, parallelization, and just-in-time compilation significantly accelerates training across all tasks, bridging the gap between the Robotarium and modern MARL algorithms.
7 Benchmarking Learning Algorithms
7.1 Experiment setup
We demonstrate the benchmarking and sim2real evaluation capabilities of JaxRobotarium by training 5 policies in parallel for all tasks using the following SOTA MARL approaches:
PQN [34]: a recent multi-agent Q-Learning approach designed to perform Q-updates efficiently by collecting non-correlated transitions in parallel across vectorized environments (we use 256), eliminating the need for a replay buffer.
QMIX [35]: an established multi-agent Q-learning algorithm where decentralized Q-networks predict agent Q-values aggregated into a global Q-value by a hypernetwork-generated centralized critic conditioned on the global state.
MAPPO [36]: an established multi-agent policy-gradient algorithm that learns decentralized action policies, regularized by a centralized critic conditioned on the global state.
IPPO [36]: a multi-agent policy-gradient algorithm that learns decentralized action policies. Unlike MAPPO, IPPO uses a decentralized critic conditioned on the agent’s local observation.
See Appendix A for further training details. We report mean episodic reward, collisions, and a task-performance metric with associated standard errors. Metrics are averaged across the 5 trained policies for 100 episodes each. For real-world experiments, we deploy the best-performing policy on the Robotarium hardware testbed and report metrics averaged across 5 unique instantiations, for a total of 160 experiments.
7.1.1 Results
| Scenario | Metric | Training-Seed Benchmark | |||
| PQN | QMIX | MAPPO | IPPO | ||
| Arctic | Reward | -3.6 0.04 | -3.4 0.02 | -3.5 0.04 | -3.5 0.04 |
| Transport | Success Rate | 1.0 0.00 | 1.0 0.00 | 1.0 0.00 | 1.0 0.00 |
| Collisions | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | |
| Discovery | Reward | 26.6 0.20 | 29.8 0.14 | 24.1 0.32 | 30.0 0.15 |
| Landmarks Tagged | 5.1 0.04 | 5.6 0.03 | 4.5 0.06 | 5.7 0.03 | |
| Collisions | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | |
| Material | Reward | -5.0 0.06 | 4.6 0.07 | 1.3 0.08 | 1.3 0.09 |
| Transport | Material Left | 65.1 0.75 | 4.2 0.31 | 15.7 0.72 | 17.1 0.71 |
| Collisions | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | |
| Warehouse | Reward | 8.4 0.20 | 6.1 0.16 | 3.9 0.02 | 8.0 0.17 |
| Deliveries | 1.7 0.06 | 0.9 0.05 | 0.0 0.00 | 1.7 0.05 | |
| Collisions | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | |
| Navigation | Reward | -18.3 0.39 | -17.0 0.17 | -13.7 0.00 | -12.7 0.16 |
| (MAPF) | Success Rate | 0.6 0.39 | 0.8 0.17 | 1.0 0.00 | 1.0 0.16 |
| Collisions | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | |
| Foraging | Reward | 1.8 0.12 | 4.5 0.13 | 0.3 0.06 | 7.1 0.21 |
| Resources Foraged | 0.3 0.02 | 0.8 0.03 | 0.1 0.01 | 1.3 0.01 | |
| Collisions | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | |
| Predator | Reward | 58.8 2.45 | 6.3 0.65 | 11.0 1.49 | 46.9 3.15 |
| Prey | Tags | 5.9 0.25 | 0.6 0.06 | 1.1 0.15 | 4.7 0.31 |
| Collisions | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | |
| Continuous | Reward | 5.9 0.78 | 3.6 0.66 | 0.0 0.59 | 1.1 0.53 |
| RWARE | Shelf Dropoffs | 5.9 0.02 | 3.6 0.02 | 0.0 0.00 | 1.1 0.01 |
| Collisions | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | |
Learning: In general, we do not find any single algorithm to consistently outperform all others across our 8 scenarios. Consistent with recent findings [34], we find that Q-Learning with PQN performs comparably or even outperforms Q-Learning with QMIX, with Material Transport being a notable exception (see Table 3). Given that PQN achieves this performance without maintaining a replay buffer, these results support PQN as a promising alternative to established Q-Learning methods for multi-robot settings, in which maintaining a large replay buffer could be intractable due to high-dimensional state spaces and combinatorial space complexity.
Further, we note that independent learning (IPPO) is surprisingly stronger in our multi-robot coordination scenarios compared to its centralized counterpart (MAPPO). We believe this is likely due to the limited partial observability in the configured benchmark scenarios (e.g., robots can generally observe their neighbors except in Navigation). This suggests that in settings without significant partial observability, independent learning is a simpler and effective alternative to centralized critic approaches, in line with prior results comparing MAPPO and IPPO on StarCraft tasks [36].
| Scenario | Metric | Simulated Experiments | Real-World Experiments | ||||||
| PQN | QMIX | MAPPO | IPPO | PQN | QMIX | MAPPO | IPPO | ||
| Arctic | Reward | -3.5 0.22 | -3.5 0.18 | -3.7 0.22 | -3.4 0.18 | -2.6 0.13 | -3.0 0.13 | -3.8 1.16 | -2.5 0.09 |
| Transport | Success Rate | 1.0 0.00 | 1.0 0.00 | 1.0 0.00 | 1.0 0.00 | 1.0 0.00 | 1.0 0.00 | 0.8 0.18 | 1.0 0.00 |
| Collisions | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | |
| Discovery | Reward | 26.1 1.88 | 25.2 1.16 | 23.6 2.68 | 29.3 1.25 | 24.7 1.07 | 26.1 1.48 | 23.9 2.55 | 29.3 1.25 |
| Landmarks Tagged | 5.0 0.40 | 4.8 0.18 | 4.6 0.54 | 5.6 0.22 | 4.6 0.22 | 5.0 0.27 | 4.6 0.45 | 5.6 0.22 | |
| Collisions | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | |
| Material | Reward | -6.1 0.04 | 3.8 0.49 | 0.4 0.85 | 0.4 0.89 | -6.1 0.04 | 4.2 0.72 | -0.3 0.89 | 1.6 0.67 |
| Transport | Material Left | 68.3 2.64 | 10.9 2.28 | 37.2 9.88 | 26.7 6.44 | 66.3 2.64 | 5.5 2.64 | 35.7 10.96 | 12.4 3.22 |
| Collisions | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | |
| Warehouse | Reward | 6.8 0.58 | 10.0 1.48 | 4.0 0.00 | 5.6 1.25 | 8.2 1.30 | 9.2 1.12 | 3.8 0.18 | 4.6 1.03 |
| Deliveries | 1.2 0.18 | 2.2 0.45 | 0.0 0.00 | 1.2 0.31 | 1.6 0.36 | 1.8 0.31 | 0.0 0.00 | 0.8 0.31 | |
| Collisions | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | |
| Navigation | Reward | -19.5 2.33 | -14.8 2.06 | -14.3 2.64 | -12.8 2.10 | -17.8 2.06 | -14.7 2.01 | -13.7 2.59 | -12.2 1.92 |
| (MAPF) | Success Rate | 0.0 0.00 | 1.0 0.00 | 1.0 0.00 | 1.0 0.00 | 0.2 0.18 | 1.0 0.00 | 1.0 0.00 | 1.0 0.00 |
| Collisions | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | |
| Foraging | Reward | 6.2 1.07 | 6.2 1.07 | 1.0 0.89 | 6.6 2.41 | 5.2 0.18 | 5.2 0.18 | 2.0 1.07 | 11.0 0.00 |
| Resources Foraged | 1.2 0.18 | 1.2 0.18 | 0.2 0.18 | 1.2 0.45 | 1.0 0.0 | 1.0 0.00 | 0.4 0.22 | 2.0 0.00 | |
| Collisions | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | |
| Predator | Reward | 44.0 17.13 | 4.0 2.19 | 14.0 8.77 | 52.0 16.37 | 14.0 5.37 | 36.0 16.64 | 6.0 3.58 | 28.1 13.64 |
| Prey | Tags | 4.4 1.70 | 0.4 0.22 | 1.4 0.89 | 5.2 1.65 | 1.4 0.54 | 3.6 1.65 | 0.6 0.36 | 3.8 1.21 |
| Collisions | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | |
| Continuous | Reward | 1.0 0.67 | 0.8 0.54 | 0.0 0.00 | 0.0 0.00 | 0.8 0.31 | 6.2 2.68 | 0.0 0.00 | 0.0 0.00 |
| RWARE | Shelf Dropoffs | 1.0 0.67 | 0.8 0.54 | 0.0 0.00 | 0.0 0.00 | 0.8 0.31 | 6.2 2.68 | 0.0 0.00 | 0.0 0.00 |
| Collisions | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | |
Sim2Real: As one would expect, we found reasonable gaps between the simulated and real world runs for most tasks across methods (see Table 4). Some of this gap could be attributed to only having limited real-world runs and the stochasticity of the real world compared to the deterministic simulator. However, two of our new tasks (Predator Prey and Continuous-RWARE) resulted in surprisingly large sim2real performance gaps, particularly for QMIX and PQN. Our qualitative analysis of the rollouts revealed that these tasks contain inflection points that can exacerbate modest differences into large differences in the environment and robot behavior over time. For instance, small positional changes to the robots in Predator-Prey can sometimes elicit qualitatively different prey behaviors which cascade into dramatic differences between the simulated and real-world rollouts. In contrast, existing MARBLER tasks are less sensitive to such variations, as robots mostly interact with a static environment. The observed gaps in the new tasks underscore the importance of hardware evaluation, especially in multi-robot scenarios where task outcomes are closely tied to interaction dynamics. The ability to accurately simulate single robots does not necessarily imply accurate multi-robot simulations.
| Scenario | Metric | Simulated Experiments | Real-World Experiments | ||
| PQN | QMIX | PQN | QMIX | ||
| Predator Prey | Returns | 30.0 6.33 | 3.3 1.21 | 16.0 5.06 | 26.0 11.33 |
| Tags | 3.0 0.62 | 0.3 0.13 | 1.6 0.52 | 2.6 1.14 | |
| Collisions | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | |
| Predator Prey | Returns | 66.7 3.23 | 16.7 6.09 | 46.7 9.66 | 18.0 4.60 |
| w/ DR | Tags | 6.7 0.34 | 1.7 0.62 | 4.7 0.96 | 1.8 0.46 |
| Collisions | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | 0.0 0.00 | |
Domain Randomization: Inspired by the above observations, we applied simple domain randomization through action noise to learn a more robust policy. We found that training with action noise can reduce the sim2real gap in Predator Prey for both QMIX and PQN and improve task performance in simulation (see Table 5).
8 Conclusion
We contribute JaxRobotarium, an accelerated benchmark platform for MRRL with free, publicly available, and standardized sim2real evaluation through the Robotarium hardware test bed. Our benchmark experiments demonstrate up to a 150x speed up in simulation, and a 20x speed up in training multi-robot policies compared to baseline. We demonstrate the utility of our platform by evaluating 4 MARL algorithms on 8 multi-robot coordination scenarios, with a large-scale sim2real evaluation consisting of over 200 real world deployments. We hope JaxRobotarium will be a valuable resource for accelerating the development, benchmarking, and sim2real evaluation of MRRL algorithms and architectures.
9 Limitations
We design our platform for sim2real evaluation through the Robotarium testbed, so we only simulate the Robotarium GRITSBots and we do not simulate observations from more complex perception modules, such as images or LiDAR scans.
The decision to utilize the Robotarium introduces additional limitations. Since the robots are homogeneous, heterogeneity has to be artificially injected, such as limiting a robot’s velocity or waypoint step size, or through introducing scenario-specific heterogeneity that is recognized by the scenario’s step function, such as obeying the foraging level of robots when deciding if a resource is successfully foraged. Furthermore, the Robotarium hardware test bed does not support Jax libraries, so the quadratic problem solver used for barrier certificates in deployment and the solver used in Jax simulation (chosen differently for compatibility with jit and vmap) are not aligned. We observe that underlying differences in the solvers can result in different proposed solutions to the same quadratic program.
We note that our platform cannot be considered a replacement for training in high-fidelity robotic simulation environments such as IsaacGym [16] and IsaacLab [17]. Rather, JaxRobotarium is designed to be a turn-key platform for simulation-training-deployment tailored for MRRL research, offering native integrations (e.g., JaxMARL), standardized benchmarks (e.g., RWARE adaptations), and a ready-to-use zero-cost sim2real pipeline.
Finally, in our training evaluations, we only use an RNN-based policy architecture since it is widely implemented to overcome the challenges of partial observability and long-horizon reasoning [35, 36, 11, 37, 12], and do not evaluate on MLP architectures [14, 15], GNN architectures [38, 39, 40, 41], or more recent Hyper-Network architectures [42, 43]. Additionally, we train with limited partial observability in most scenarios, see Appendix B for exact details. Future work could benchmark these policy architectures within JaxRobotarium with stronger partial observability.
Appendix A Training Details
A.1 Policy Architecture
We use an RNN-based policy architecture across all of our training experiments. The specific implementations are all adopted from JaxMARL’s baseline implementations [11] and use a Gated Recurrent Unit (GRU) [44].
PQN: A 2 layer MLP with Layer Normalization before ReLU activations, followed by a GRU, followed by a linear layer.
QMIX: A 1 layer MLP with ReLU activation, followed by a GRU, followed by a linear layer.
MAPPO: A single layer MLP with ReLU activation, followed by a GRU, followed by a single layer MLP with ReLU activation, followed by a linear layer.
IPPO: A single layer MLP with ReLU activation, followed by a GRU, followed by a single layer MLP with ReLU activation, followed by a linear layer.
For all algorithms, we train 5 policies in parallel across 5 unique seeds. Exact widths for the policies’ hidden layers were chosen by ablating over the widths 128, 256, and 512 trained on a single seed. The best-performing width was then chosen for the final training runs. We report all policy widths in Table 6.
| Scenario | PQN | QMIX | MAPPO | IPPO |
| Arctic | 256 | 256 | 128 | 128 |
| Transport | ||||
| Discovery | 256 | 256 | 128 | 256 |
| Material | 512 | 128 | 128 | 512 |
| Transport | ||||
| Warehouse | 256 | 512 | 128 | 256 |
| Navigation | 256 | 512 | 128 | 512 |
| Foraging | 128 | 128 | 512 | 256 |
| Predator | 128 | 512 | 256 | 512 |
| Prey | ||||
| Continuous | 256 | 256 | 128 | 256 |
| RWARE |
A.2 PQN
| Parameter | Value |
| NUM_ENVS | 256 |
| NORM_INPUTS | False |
| NORM_TYPE | layer_norm |
| EPS_START | 1.0 |
| EPS_FINISH | 0.1 |
| EPS_DECAY | 0.1 |
| MAX_GRAD_NORM | 1 |
| NUM_MINIBATCHES | 16 |
| NUM_EPOCHS | 4 |
| LR | 0.0005 |
| LR_LINEAR_DECAY | False |
| GAMMA | 0.9 |
| LAMBDA | 0.3 |
We train with the JaxMARL tuned hyperparameters listed in Table 7.
A.3 QMIX
| Parameter | Value |
| NUM_ENVS | 8 |
| BUFFER_SIZE | 5000 |
| BUFFER_BATCH_SIZE | 32 |
| MIXER_EMBEDDING_DIM | 32 |
| MIXER_HYPERNET_HIDDEN_DIM | 128 |
| MIXER_INIT_SCALE | 0.001 |
| EPS_START | 1.0 |
| EPS_FINISH | 0.05 |
| EPS_DECAY | 0.1 |
| MAX_GRAD_NORM | 25 |
| TARGET_UPDATE_INTERVAL | 200 |
| TAU | 1.0 |
| NUM_EPOCHS | 1 |
| LR | 0.005 |
| LEARNING_STARTS | 10000 |
| LR_LINEAR_DECAY | True |
| GAMMA | 0.9 |
We train with the JaxMARL tuned hyperparameters listed in Table 8.
A.4 MAPPO
| Parameter | Value |
| LR | 0.002 |
| NUM_ENVS | 16 |
| HIDDEN_SIZE | 128 |
| UPDATE_EPOCHS | 4 |
| NUM_MINIBATCHES | 4 |
| GAMMA | 0.99 |
| GAE_LAMBDA | 0.95 |
| CLIP_EPS | 0.2 |
| SCALE_CLIP_EPS | False |
| ENT_COEF | 0.01 |
| VF_COEF | 0.5 |
| MAX_GRAD_NORM | 0.5 |
| ANNEAL_LR | True |
We train with the JaxMARL tuned hyperparameters listed in Table 9.
A.5 IPPO
| Parameter | Value |
| LR | 0.0005 |
| NUM_ENVS | 16 |
| UPDATE_EPOCHS | 4 |
| NUM_MINIBATCHES | 2 |
| GAMMA | 0.99 |
| GAE_LAMBDA | 0.95 |
| CLIP_EPS | 0.3 |
| SCALE_CLIP_EPS | False |
| ENT_COEF | 0.01 |
| VF_COEF | 1.0 |
| MAX_GRAD_NORM | 0.5 |
| ANNEAL_LR | True |
We train with the JaxMARL tuned hyperparameters listed in Table 10.
Appendix B Scenario Details
As a general note for all scenarios, if heterogeneity awareness is configured, robots will additionally observe the configured heterogeneity representation. In our training experiments, the action space for all robots is waypoints that are followed by a unicycle position controller with barrier functions enabled for collision avoidance.
B.1 Arctic Transport
B.1.1 Description
Adopted from MARBLER, in this scenario, the ice and water robots must cross the terrain consisting of ice tiles (light blue), water tiles (dark blue), and ground tiles (white) to reach the goal zone (green). Ice and water robots move at the same speed across ground tiles, slow over tiles not corresponding to their type, and fast over tiles corresponding to their types. Ice and water robots observe their position, the position of other robots, and the current tile type they are occupying. Drone robots observe a 3x3 grid centered around their current position, and move at a uniform speed across all tiles. Their observations are directly communicated to the ice and water robots, such that they can help guide the ice and water robots. The team is rewarded for minimizing the distance of the ice and water robots to the goal zone, and penalized for every step taken where both robots are not in the goal zone.
B.1.2 Experiment Parameters
Number of Robots: 4
Max steps: 100
Training timesteps: 3e6
Controller
Controller: unicycle Position
Barrier Fn: enabled
Normal Step: 0.2
Fast Step: 0.3
Slow Step: 0.1
Heterogeneity
Type: class
Obs Type: class (robots observe their own class)
Values: [[1, 0, 0], [1, 0, 0], [0, 1, 0], [0, 0, 1]] (each entry indicates the type [drone, water, ice])
Distance Reward: -0.05 summed shortest distances of ice and water robots to the goal zone
Step Reward: -0.10 0 if all robots are on their goal, otherwise 1
Violation Reward: 0 number of collisions
B.2 Discovery
B.2.1 Description
Renamed from Predator-Capture-Prey in MARBLER to better reflect the scenario objective, in this scenario, N sensing robots must collaborate with M tagging robots to tag K landmarks (black) in the environment. Sensing robots and tagging robots have configurable sensing radii. Robots observe their own location and the location of other robots. Initially, robot observations also contain a position outside the map for all landmarks, indicating they have not been sensed. When a landmark is within range of a sensing robot, it is considered sensed, the black marker is updated to green, and the location becomes available within the observation. When a landmark is within range of a tagging robot, the landmark is considered tagged and disappears from the visualization. The team is rewarded for landmarks sensed and tagged, and penalized for every step taken where all landmarks are not tagged. This scenario can scale to an arbitrary number of robots and landmarks.
B.2.2 Experiment Parameters
Number of Robots: 4
Sensing Robots: 2
Tagging Robots: 2
Max Steps: 100
Training timesteps: 5e6
Controller
Controller: unicycle position
Barrier Fn: enabled
Heterogeneity
Type: capability set
Obs type: full capability set (robots observe the team’s capability values)
Values: [[0.45, 0], [0.45, 0], [0, 0.25], [0, 0.25]] (each entry is [sensing radius, tag radius])
Sensing Reward: 1 landmarks sensed
Tagging Reward: 5 landmarks tagged
Step Reward: -0.05 0 if all landmarks tagged, otherwise 1
Violation Reward: 0 number of collisions
B.3 Foraging
B.3.1 Description
Inspired by grid-world level-based foraging in [31], in this scenario, N robots with varying levels must collaborate to forage M resources (black markers) within the environment. Labels indicate the levels of robots and resources. Robots observe their position, the position of other robots, the location of resources, and their corresponding levels. For a resource to be considered foraged, it must be within the foraging radius of M robots, where the summed level of the M robots is greater than or equal to the level of the resource. When resources are foraged, they disappear from the visualization to indicate they have been collected. The team is rewarded for each resource foraged, scaled by the level of that resource. This scenario can scale to an arbitrary number of robots and resources.
B.3.2 Experiment Parameters
Number of Robots: 3
Number of Resources: 2
Max steps: 100
Training timesteps: 5e6
Controller
Controller: unicycle position
Barrier Fn: enabled
Heterogeneity
Type: capability set
Obs type: full capability set (robots observe the team’s capability values)
Values: [[1], [2], [3]] (each entry is the robot’s foraging level)
Foraging Reward: 1 level of resources foraged
Violation Reward: 0 number of collisions
B.4 Material Transport
B.4.1 Description
Adopted from MARBLER, in this scenario, N robots with varying speeds and carrying capacities must collaborate to unload two loading zones (green circle and green rectangle) into the drop-off zone (purple rectangle). Both zones are initialized with an amount of material sampled from a zone-specific distribution. We assume both distributions are Gaussian with pre-specified means and variances. Robots observe their position, the position of other robots, and the remaining material to unload in both zones. Robots can only load material by entering the loading zones, can only carry the amount of material specified to their carrying capacity, and must drop off their loaded material into the drop-off zone before loading more material. The team is rewarded for the amount of material loaded and the amount of material dropped off, and it is penalized for each step taken where there is material remaining in the loading zones. This scenario can scale to an arbitrary number of robots.
B.4.2 Experiment Parameters
Number of Robots: 4
Max steps: 70
Training timesteps: 5e6
Controller
Controller: unicycle position
Barrier Fn: enabled
Heterogeneity
Type: capability set
Obs type: full capability set (robots observe the team’s capability values)
Values: [[0.45, 5], [0.45, 5], [0.15, 15], [0.15, 15]] (each entry is [step size, carrying capacity])
Gaussian distribution parameters of materials in circular loading zone
Mean: 75
Variance: 10
Gaussian distribution parameters of materials in rectangular loading zone
Mean: 15
Variance: 4
Load Reward: 0.25 material loaded by robots
Drop Off Reward: 0.75 material dropped off by robots
Step Reward: -0.1 0 if no material remaining in loading zones, otherwise 1
Violation Reward: 0 number of collisions
B.5 Navigation (MAPF)
B.5.1 Description
MAPF is a classic problem in multi-robot/multi-agent systems [45, 46]. In this scenario, N robots must navigate to individually specified goals (black markers). Robots observe their position and the vector to their goal location. The team is rewarded for robots minimizing their distance to their respective goals. This scenario can scale to an arbitrary number of robots.
B.5.2 Experiment Parameters
Number of Robots: 3
Max steps: 100
Training timesteps: 3e6
Controller
Controller: unicycle position
Barrier Fn: enabled
Distance Reward: -1 euclidean distance of robots to their respective goals
Violation Reward: 0 number of collisions
B.6 Predator Prey
B.6.1 Description
Based on a variant of the classic predator-prey scenario introduced in [47], in this scenario, N tagging robots must collaborate to tag a more agile prey (green marker). The prey is controlled by the heuristic proposed in [33], where at each step the prey maximizes its distance from the closest tagging robot. Robots observe their position, the position of other robots, and the position of the prey. Robots tag the prey by getting the prey within their tagging radius, and the prey marker will briefly turn red to indicate a successful tag. The team is rewarded for each successful tag on the prey. This scenario can scale to an arbitrary number of robots.
B.6.2 Experiment Parameters
Number of Robots: 3
Max steps: 100
Training timesteps: 10e6
Controller
Controller: unicycle position
Barrier Fn: enabled
Tag Reward: 10 1 if a tag occurred, 0 otherwise
Violation Reward: 0 number of collisions
B.7 Continuous-RWARE
B.7.1 Description
Inspired by grid-world RWARE [32], in this scenario, robots must collaborate to deliver the requested shelves (indicated by the top left text) to the dropoff zone (purple), where the request is updated upon each successful shelf delivery. Shelves are individually labeled and begin in staging zones (gray) arranged in a grid. Robots observe their position, the position of other robots, the position of each staging zone and the shelf occupying it, and the requested set of shelves. Robots can pickup or return shelves by entering the gray staging zone. If a robot is not carrying a shelf, it can freely move underneath shelves and through staging zones. Once a robot is carrying a shelf, it cannot move through other shelves, and shelves can only be returned at an unoccupied staging zone. A robot successfully completes a drop-off when it enters the purple zone carrying a shelf that is currently requested, at which point the corresponding entry in the request set is randomly updated to any non-requested shelf. Robots are sparsely rewarded for each successful drop-off. This scenario scales to an arbitrary number of robots and shelves.
B.7.2 Experiment Parameters
Number of Robots: 3
Number of Shelves: 6
Max steps: 100
Training timesteps: 20e6
Controller
Controller: unicycle position
Barrier Fn: enabled
Drop Off Reward: 1 number of successful shelf drop-offs
Violation Reward: 0 number of collisions
B.8 Warehouse
B.8.1 Description
Adopted from MARBLER, in this scenario, N red robots and M green robots must navigate to the zone with matched color on the right to pick up packages, and then deliver them to the zone with corresponding colors on the left. Robots observe their position and the position of other robots. The team is rewarded for packages loaded and packages delivered. This scenario scales to an arbitrary number of robots.
B.8.2 Experiment Parameters
Number of Robots: 4
Max steps: 70
Training timesteps: 3e6
Controller
Controller: unicycle position
Barrier Fn: enabled
Heterogeneity
Type: class
Obs type: class (robots observe their own class)
Values: [[1, 0], [1, 0], [0, 1], [0, 1]] (each entry indicates the type [green, red])
Load Reward: 1 packages loaded
Delivery Reward: 3 packages delivered
Violation Reward: 0 number of collisions